OpenStudy (anonymous):

Find the exact value of tan5^-1 (1)

OpenStudy (jdoe0001):

\(\bf tan^{-1}(1)\quad ?\)

OpenStudy (anonymous):

nah theres a 5 before the -1

OpenStudy (anonymous):

thats why im confused

OpenStudy (jdoe0001):

can you post a picture of it?

OpenStudy (anonymous):

makes no sense to me

OpenStudy (jdoe0001):

tis a typo, is noticeable due to the choices given

OpenStudy (jdoe0001):

keep in mind that the ^ is next to the "5" on the keyboard =)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

so just 7pi/4 then?

OpenStudy (jdoe0001):

\(\bf tan^{-1}(1)\implies tan(\theta)=1\implies tan^{-1}(tan(\theta))=tan^{-1}(1)\\ \quad \\ \theta=tan^{-1}(1)\) yes, that one would work keep in mind that for the inverse tangent function, its domain is \(\bf \left(\cfrac{\pi}{2},-\cfrac{\pi}{2}\right)\) so that would fall under that range

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw